Calculate the number of gallons of paint required to paint the outside walls of a barn that measures 50 feet by 100 feet. Each wall is 12 feet high and each gallon of paint covers 600 square feet.
The answer I came up with is 8.3 which is wrong.

Found 2 solutions by checkley77, Alan3354:
Answer by checkley77(12844)

(2*50+2*100)*12
(100+200)12
300*12=3,600 ft^2
3,600/600=6 gallons of paint will be required.

Answer by Alan3354(69443)

Calculate the number of gallons of paint required to paint the outside walls of a barn that measures 50 feet by 100 feet. Each wall is 12 feet high and each gallon of paint covers 600 square feet.
The answer I came up with is 8.3 which is wrong.
---------------------------
I think you used 50*100, which is the area of the floor.
--------------------------------
Find the area of the walls. Two are 12 by 50 = 600 sf each. The other 2 are 12 by 100 = 1200 sf each.
The total is 2*600 + 2*1200 = 3600 sq feet
3600/600 = 6 gallons.